OpenStudy (anonymous):

Integral of x^2/(x^3+1) which is the general solution to start off with?

OpenStudy (dls):

Let x^3+1=u Differentiate both sides to get : 3x^2 = du/dx dx=du/3x^2 Now you have to integrate simply\[\Huge \int\limits \frac{x^2 }{u } \times \frac{du}{3x^2}\]

OpenStudy (anonymous):

I started off with u = x^2 and du = 2x dv = 1/x^3+1 that gives v = kn (x^3+1) then u'v+uv'

OpenStudy (dls):

nope you can't let u=x^2

OpenStudy (anonymous):

oh okey thanks, I'll try to solve it.

OpenStudy (dls):

:)

OpenStudy (anonymous):

okey so I shorten it down first to: \[\frac{ 1 }{ 3 } \int\limits_{}^{} \frac{ x^2 }{ x^3+1 } * \frac{ 3x^2 }{ 3x^2 }\] then the remaining equation is: \[\frac{ 1 }{ 3 } \int\limits_{}^{} \frac{ x^2 }{ x^3 +1}\] Where do I go from here?

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

Okey first misstake realised is the 1/3, it should be remoived.

OpenStudy (anonymous):

\[\frac{ x^3 }{ 3 } \ln x^3+1 +c\] now how does that convert into \[\frac{ 1 }{ 3 } \ln x^3 +1+ c\]

OpenStudy (kainui):

Hmm this seems fishy to me, let's start over. \[\Large \int\limits \frac{x^2}{x^3+1}dx\] Now we should notice that the derivative of x^3 is really just x^2 times a constant. That is usually a good sign that we should pick x^3 to be used for u-substitution. Since the derivative of a constant is 0, we also pick the 1 along with it \[\Large u=x^3+1\] Now we take the derivative \[\Large \frac{du}{dx}=3x^2\]and then we rearrange it so it looks like what we have left so that we can plug it into the dx part of the integral with the x^2.\[\Large \frac{du}{3}=x^2dx\]See? Just algebra. Ok, so if we plug this in, I'll sort of show you in steps to see how I'm doing this: \[\Large \int\limits \frac{x^2}{x^3+1}dx=\int\limits (\frac{1}{x^3+1})*(x^2dx)\] Now we substitute thes things both in: \[\Large \int\limits (u)*(\frac{du}{3})=\frac{1}{3} \int\limits u du\]Now you can siply integrate this and then take your original u from earlier and plug it into the answer to get it back in terms of x.

OpenStudy (kainui):

Whoops I think I messed up right at the end there haha! \[\Large \int\limits (\frac{1}{x^3+1})*(x^2dx)=\int\limits (\frac{1}{u})*(\frac{du}{3})\]

OpenStudy (anonymous):

okey I solved it, with your help, sometimes I make it harder than it needs to be. \[\frac{ 1 }{ 3 } \ln x^3+1 +c\]

OpenStudy (zarkon):

\[\Large \int\frac{1}{u}du=\ln|u|+c\]

OpenStudy (anonymous):

yes, I made the misstake by putting in the numbers in du and u and started dividing and so on. @Kainui your explanation was very clear, thanks.

OpenStudy (larseighner):

Thanks to Leibniz, infinitesimals do work like algebra in operations like these. I think it is a stretch to say these operations are algebra.

OpenStudy (kainui):

I think it's a stretch to say it's not algebra.

OpenStudy (anonymous):

Okey I'll close this now, and continue with a simular integral and hopefully be able to solve, thanx to evry1 for taking their time!:)